# Tool Zero: Training Tool-Augmented LLMs via Pure RL from Scratch

### v1

@@ Abstract
Yirong Zeng1, Xiao Dingthanks: Corresponding author. Email: xding@ir.hit.edu.cn 1, Yutai Hou2, Yuxian Wang2, Li Du4, Juyi Dai1, Qiuyang Ding2, Duyu Tang3, Dandan Tu3, Weiwen Liu3, Bing Qin1, Ting Liu1, 1Harbin Institute of Technology SCIR Lab, 2Huawei Technologies Co., Ltd, 3Shanghai Jiao Tong University, 4Beijing Academy of Artificial Intelligence
@@ 1 Introduction
Integrating LLMs with external tools has emerged as a pivotal advancement, significantly enhances their ability to address complex tasks (Qu et al., 2025; Wang et al., 2024). It opens up many practical uses across different fields. For example, it supports the automation of reasoning tasks (Jin et al., 2025; Manduzio et al., 2024), and enables Agent applications (Gunter et al., 2024; Chen et al., 2024). A tool-augmented model can respond to a user’s query by invoking and executing external tools. In this paper, tools are used interchangeably with APIs, functions, and plugins.
@@
Current approaches to enhance tool-use capability involve synthesizing extensive tool-use trajectories with advanced language models, followed by SFT on the generated data (Liu et al., 2024a; Lin et al., 2024; Chen et al., 2025). Under this paradigm, models achieves satisfactory performance in the scenarios that have the same distribution as the training data (Yan et al., 2024). However, these SFT-trained models primarily engage in imitating surface-level patterns rather than internalizing the reasoning process, tend to memorize the training trajectories rather than developing robust, intrinsic reasoning capabilities (Chen et al., 2025). Consequently, they exhibit limited generalization ability when applied to unseen scenarios, as elaborated in preliminary study in Section 3.
@@
Recent studies have demonstrated that simple rule-based R1-style RL (DeepSeek-AI, 2025), even without SFT, can significantly enhance LLMs’ complex reasoning capabilities (Zeng et al., 2025a; Lu et al., 2025; Shen et al., 2025). This paradigm inspires us to extend pure RL to the tool learning domain, aiming to address the generalization limitations by eliciting models’ intrinsic reasoning abilities. To this end, we propose a dynamic generalization-guided reward design for rule-based RL. This approach employs a progressive reward strategy: it first promotes early-stage exploratory behavior to cultivate intrinsic reasoning, then refines these capabilities into tool-use patterns focused on final-task precision. This design effectively resolves the exploration-exploitation dilemma in open-domain tool learning, bridging the gap between reasoning generalization and task-specific tool use.
@@ 2.1 Tool Learning
Enhancing LLMs with external tools has emerged as a pivotal direction for addressing complex tasks in open domains (Qu et al., 2025; Wang et al., 2024). Typical applications include integrating LLMs with search engines (Zhang et al., 2024b; Lazaridou et al., 2022; Shuster et al., 2022), calculators (Nakano et al., 2021), and Python interpreters (Wang et al., 2024; Song et al., 2024; Chen et al., 2022). A dominant paradigm for equipping LLMs with external tools is imitation learning, where language models are trained via imitation on human-labeled datasets. This framework typically involves constructing large-scale supervised tool-use datasets (Prabhakar et al., 2025; Liu et al., 2024b, a) and applying either SFT (Zhang et al., 2024a, b; Qin et al., 2023) or direct preference optimization (DPO) reinforcement learning (Zeng et al., 2025b; Yu et al., 2024), enabling models to autonomously create and invoke tools. However, this paradigm faces challenges in enabling LLMs to generalize across diverse tools with varied argument structures and domains, highlighting a critical gap in tool-agnostic generalization.
@@ 2.2 Tool-Integrated Reasoning with Reinforcement Learning
RL has gained traction as a more scalable and generalizable training paradigm. Models like R1-Zero leverage group relative policy optimization (GRPO) (Shao et al., 2024) to unlock the model’s reasoning capabilities at test time (DeepSeek-AI, 2025; Yu et al., 2025). This R1-style reasoning paradigm, marking a shift from train-time scaling to test-time scaling (Muennighoff et al., 2025; Xia et al., 2025), has demonstrated success in mathematics (Shao et al., 2024), coding (Pan and Liu, 2025).
Recent studies (Jin et al., 2025; Qian et al., 2025) have explored unlocking tool-integrated reasoning for LLMs, with works like Torl (Li et al., 2025) and ReTool (Feng et al., 2025) achieving promising performance in mathematical tasks by integrating code tools. However, their training follows the SFT-then-RL paradigm and remains constrained to single-type tool-use scenarios. In contrast, our work aims to unlock the model’s tool-agnostic (general-purpose tools) generalization capabilities via pure reinforcement learning scaled directly from Zero model.
@@ 3.1 Preliminary Study
(1) Intra-Dataset Performance. We compared two SFT-trained models: ToolACE-8B (Llama3.1-8b-inst finetuned on ToolACE (Liu et al., 2024a)) and xLAM-7B-r (Mistral-7b finetuned on xLAM (Liu et al., 2024b)), evaluated on the BECL benchmark (Yan et al., 2024) (comprising Single-turn (Non-Live, Live) and Multi-turn subsets). Notably, both training datasets use LLM-synthesized data to mimic real-world scenarios, with distributions aligned to BFCL-Live (details in Section 4.1). Results in Figure 2 show significant improvements on the Live metric (reflecting in-distribution performance), such as a improvement from 61.1 to 78.6. Conversely, exhibiting negligible gains or regressions on Non-Live and Multi-Turn subsets (e.g., 9.6 $\rightarrow$ 7.8). This suggests that SFT struggles with out-of-distribution generalization in open-domain settings. For instance, single-turn training data fails to transfer to multi-turn scenarios, and simple tool use patterns do not generalize to complex, interdependent tool chains.
(2) Cross-Dataset Performance. We extended our evaluation to diverse benchmarks (details in Appendix B), inspired by Lin et al., 2024. Notably, these benchmarks encompass varied tool-use scenarios(e.g., candidate tools, contextual domains, and invocation formats (JSON vs. Python code)). Results in Table 1 reveal inconsistent performance of existing tool-use models across benchmarks. For example, while xLAM-7B-fc achieved top performance on BFCL, it suffered significant degradation on two others, leading to the lowest overall average score. In contrast, the foundation models demonstrated more consistent cross-dataset performance. Therefore, this result highlighting a critical issue: SFT enhances in-distribution performance but weakens generalization to unseen scenarios. (e.g., novel tools, invocation formats).
@@ 4.1 Training Data Preparation
The following data are utilized in RL training. (1) ToolACE (Liu et al., 2024a): It is a general-purpose tool-use dataset, where the model learns when to invoke tools and when to respond directly, thereby enhancing decision-making in multi-step interactions. (2) xLAM (Liu et al., 2024b): This is a compositional dataset that requires one or multiple tool calls per turn. It encourages the model to reason about tool dependencies and actively plan diverse tool-calling actions. We also include an irrelevance-augmented subset [^1] https://huggingface.co/datasets/MadeAgents/xlam-irrelevance-7.5k originating from xLAM.
@@
Data Mask. Recent studies (Lin et al., 2024; Chen et al., 2025) have shown that naming preferences in tool descriptionss can significantly degrade model robustness when testing environments diverge from training conventions. To mitigate this issue, we adopt a name-masking strategy aligned with Lin et al., 2024, which masks function names (e.g., calculate_sumfunc_1) and parameter names (e.g., input_listparam_1). It redirects the model’s attention to tool descriptions and argument semantics, reduces overfitting to superficial naming patterns, thereby improving tool-agnostic reasoning in open-domain settings.
@@ 4.2 Generalization Guided Reward Design
To enhance tool generalization within the GRPO framework, we introduce a dynamic generalization-guided reward design that combines flexible exploration with structured convergence. Building on prior rule-based reward mechanisms (Qian et al., 2025; Li et al., 2025; Jin et al., 2025), our formulation decomposes the total reward $\mathcal{R}_{\text{final}}$ into two components:
@@
To further enhance the model’s tool-integrated reasoning in context, we incorporate two tool-specific feedback signals into the strict reward $\mathcal{R}_{\text{strict}}$: (1) multi-tool collaboration: +0.3+0.3 reward for reinforcing correct collaborative tool usage patterns in multi-step tool chains or dialogs. (2) parameter value error: +0.3+0.3 penalty per invalid parameter value to enforce exploring precise context grounding because it demands high-order reasoning from context (Zeng et al., 2025b; Lin et al., 2024).
@@ 4.3 RL Training with Generalization-guided Reward
To train LLMs from Zero model through scaling reinforcement learning, we employ Group Relative Policy Optimization (GRPO) (DeepSeek-AI, 2025; Shao et al., 2024), that unlocks the model’s reasoning capabilities at test time.
@@
here, $\epsilon$ is a hyperparameter. Following Yu et al., 2025, we remove the KL divergence regularization from the GRPO objective. And $A_{i}$ is the advantage computed using a group of rewards $\{r_{1},r_{2},\dots,r_{G}\}$ corresponding to the completions within each group:
@@ 5.1 Experimental Setup
Evaluation Dataset. The BFCL evaluates the LLMs ability to invoke functions in the real-world by actually triggering the API call and comparing responses, provides a comprehensive dataset comprising 4k+ instances (updating), consisting of Non-live, Live (with user-contributed complex tools avoiding contamination), Multi-turn subset. Other benchmarks, namely API-Bank (Li et al., 2023), Nexus Raven (Srinivasan et al., 2023), Tool-Alpaca (Tang et al., 2023), and Seal-Tools (Wu et al., 2024), are elaborated in Appendix B.
Baselines (1) Vanilla Model: the original model without additional training (e.g., Llama3.1-series, Qwen2.5-series). (2) SFT Models: instruct models fine-tuned on supervised data, to assess whether GRPO training outperforms standard SFT, including ToolACE-8B (trained in ToolACE), xLAM-series (trained in xLAM)(Zhang et al., 2024a), and Hammer-series (trained on xLAM with function mask to enhance generalization) (Lin et al., 2024). (3) API-based closed-source models (e.g., GPT-series, Gemini-series). (4) R1-like Model: models trained using GRPO with SFT as the RL paradigm, such as QwQ-32B (Team, 2025), Tool-N1 series (single turn tool-use models trained in mixed ToolACE and xLAM data) (Zhang et al., 2025), and ToolRL(trained in subset of mixed ToolACE and xLAM data) (Qian et al., 2025). Instead, our Tool-Zero series trained with pure RL without SFT (i.e., R1-Zero).
@@ 5.3.1 Ablation Study
Additionally, we conduct a hyperparameter ablation study on the progressive reward strategy by varying two key parameters: transition midpoint $t_{m}\in\{0,25,50,100\}$ and steepness factor $\kappa\in\{0.2,1.0\}$ (controlling reward transition slope). Results (Figure 5) show that a smaller transition midpoint ($t_{m}=25$) yields the best performance, while larger values ($t_{m}\geq 50$) lead to degradation. This aligns with prior observations (Pan et al., 2024; Zhang et al., 2025) that excessive exploration in fine-grained schemes may induce reward hacking and overfitting to superficial cues. Also, a lower steepness factor consistently outperforms, indicating that gradual reward shaping stabilizes training. These findings validate the design choices in GG-GRPO’s progressive reward mechanism.
@@ A.1 The Implementation Settings
The BFCL is an evolving benchmark. For our study, we utilized the version checked out on February 26, 2024. Other benchmarks include: (1) API-Bank with 314 tool-use dialogues and 753 API calls, evaluating known API invocation (L-1) and candidate list retrieval/calling (L-2), we report their average result in evaluation; (2) Nexus Raven API Evaluation offering 318 test examples across 65 APIs for function-calling assessment; (3) Tool-Alpaca’s 271 synthetic tool-use instances in 50 categories (100 simulated tests used); (4) Seal-Tools, a recent benchmark with 4,076 auto-generated APIs across life domains. The BFCL assesses models using Abstract Syntax Tree Evaluation and Executable Function Evaluation Accuracy, and the other benchmarks assesses models using Function and Parameter matching F1 score (Lin et al., 2024). For the evaluation, each model was administered the same system prompt as its training counterpart.
@@ Appendix C System Thinking Template
We adopt a lightweight prompting schema to elicit tool-use capabilities from the LLM, drawing inspiration from prior work DeepSeek-AI (2025); Face (2025). As illustrated in Figure 6, the template explicitly instructs the model to encapsulate intermediate reasoning within <think>...</think> tags, followed by the final answer enclosed in <answer>...</answer> tags. By allowing the model greater freedom in articulating its reasoning process, we aim to enhance generalization across diverse tool integration scenarios. Additionally, this design facilitates seamless adaptation to complex tool-augmented reasoning tasks.
@@ Appendix D Model Scale Analysis
Additionally, for the experiment on algorithmic effectiveness, the objective was to evaluate the reward function on native PPO (Song et al., 2025; Zeng et al., 2025a) (i.e., GG-PPO) and GRPO algorithms (i.e., GG-GRPO), as shown in Table 9. The results revealed that our reward module with GRPO outperformed PPO, demonstrating robust and consistent gains. This highlights GRPO’s superior adaptability to the reward framework for tool learning, which can be attributed to the reward function being designed for rule-based rewards in algorithms like GRPO (unlike PPO, which is natively for model-based rewards) and recent studies showing GRPO’s advantage in achieving "aha moments" compared to PPO-family algorithms (Qian et al., 2025; DeepSeek-AI, 2025; Muennighoff et al., 2025).

### v2

@@ Abstract
Yirong Zeng1, Xiao Dingthanks: Corresponding author. Email: xding@ir.hit.edu.cn 1, Yutai Hou2, Yuxian Wang2, Li Du4, Juyi Dai1, Qiuyang Ding2, Duyu Tang2, Dandan Tu2, Weiwen Liu3, Bing Qin1, Ting Liu1, 1Harbin Institute of Technology SCIR Lab, 2Huawei Technologies Co., Ltd, 3Shanghai Jiao Tong University, 4Beijing Academy of Artificial Intelligence
@@ 1 Introduction
Integrating LLMs with external tools has emerged as a pivotal advancement, significantly enhances their ability to address complex tasks (qu2025tool; wang2024executable). It opens up many practical uses across different fields. For example, it supports the automation of reasoning tasks (jin2025search; manduzio2024improving), and enables Agent applications (gunter2024apple; chen2024octopus). A tool-augmented model can respond to a user’s query by invoking and executing external tools. In this paper, tools are used interchangeably with APIs, functions, and plugins.
@@
Current approaches to enhance tool-use capability involve synthesizing extensive tool-use trajectories with advanced language models, followed by SFT on the generated data (liu2024toolace; lin2024hammer; chen2025acebench). Under this paradigm, models achieves satisfactory performance in the scenarios that have the same distribution as the training data (bfclv3). However, these SFT-trained models primarily engage in imitating surface-level patterns rather than internalizing the reasoning process, tend to memorize the training trajectories rather than developing robust, intrinsic reasoning capabilities (chen2025acebench). Consequently, they exhibit limited generalization ability when applied to unseen scenarios, as elaborated in preliminary study in Section 3.
@@
Recent studies have demonstrated that simple rule-based R1-style RL (deepseekai2025r1), even without SFT, can significantly enhance LLMs’ complex reasoning capabilities (zeng2025simplerl; lu2025ui; shen2025vlm). This paradigm inspires us to extend pure RL to the tool learning domain, aiming to address the generalization limitations by eliciting models’ intrinsic reasoning abilities. To this end, we propose a dynamic generalization-guided reward design for rule-based RL. This approach employs a progressive reward strategy: it first promotes early-stage exploratory behavior to cultivate intrinsic reasoning, then refines these capabilities into tool-use patterns focused on final-task precision. This design effectively resolves the exploration-exploitation dilemma in open-domain tool learning, bridging the gap between reasoning generalization and task-specific tool use.
@@ 2.1 Tool Learning
Enhancing LLMs with external tools has emerged as a pivotal direction for addressing complex tasks in open domains (qu2025tool; wang2024executable). Typical applications include integrating LLMs with search engines (zhang2024ecoact; lazaridou2022internet; shuster2022blenderbot), calculators (nakano2021webgpt), and Python interpreters (wang2024executable; song2024adaptive; chen2022program). A dominant paradigm for equipping LLMs with external tools is imitation learning, where language models are trained via imitation on human-labeled datasets. This framework typically involves constructing large-scale supervised tool-use datasets (prabhakar2025apigen; liu2024apigen; liu2024toolace) and applying either SFT (zhang2024xlam; zhang2024ecoact; toolllm) or direct preference optimization (DPO) reinforcement learning (zeng2025boosting; yu2024steptool), enabling models to autonomously create and invoke tools. However, this paradigm faces challenges in enabling LLMs to generalize across diverse tools with varied argument structures and domains, highlighting a critical gap in tool-agnostic generalization.
@@ 2.2 Tool-Integrated Reasoning with Reinforcement Learning
RL has gained traction as a more scalable and generalizable training paradigm. Models like R1-Zero leverage group relative policy optimization (GRPO) (shao2024deepseekmath) to unlock the model’s reasoning capabilities at test time (deepseekai2025r1; yu2025dapo). This R1-style reasoning paradigm, marking a shift from train-time scaling to test-time scaling (muennighoff2025s1; xia2025generative), has demonstrated success in mathematics (shao2024deepseekmath), coding (pan2025metaspatial).
Recent studies (jin2025search; qian2025toolrl) have explored unlocking tool-integrated reasoning for LLMs, with works like Torl (li2025torl) and ReTool (feng2025retool) achieving promising performance in mathematical tasks by integrating code tools. However, their training follows the SFT-then-RL paradigm and remains constrained to single-type tool-use scenarios. In contrast, our work aims to unlock the model’s tool-agnostic (general-purpose tools) generalization capabilities via pure reinforcement learning scaled directly from Zero model.
@@ 3.1 Preliminary Study
(1) Intra-Dataset Performance. We compared two SFT-trained models: ToolACE-8B (Llama3.1-8b-inst finetuned on ToolACE (liu2024toolace)) and xLAM-7B-r (Mistral-7b finetuned on xLAM (liu2024apigen)), evaluated on the BECL benchmark (bfclv3) (comprising Single-turn (Non-Live, Live) and Multi-turn subsets). Notably, both training datasets use LLM-synthesized data to mimic real-world scenarios, with distributions aligned to BFCL-Live (details in Section 4.1). Results in Figure 2 show significant improvements on the Live metric (reflecting in-distribution performance), such as a improvement from 61.1 to 78.6. Conversely, exhibiting negligible gains or regressions on Non-Live and Multi-Turn subsets (e.g., 9.6 $\rightarrow$ 7.8). This suggests that SFT struggles with out-of-distribution generalization in open-domain settings. For instance, single-turn training data fails to transfer to multi-turn scenarios, and simple tool use patterns do not generalize to complex, interdependent tool chains.
(2) Cross-Dataset Performance. We extended our evaluation to diverse benchmarks (details in Appendix B), inspired by lin2024hammer. Notably, these benchmarks encompass varied tool-use scenarios(e.g., candidate tools, contextual domains, and invocation formats (JSON vs. Python code)). Results in Table 1 reveal inconsistent performance of existing tool-use models across benchmarks. For example, while xLAM-7B-fc achieved top performance on BFCL, it suffered significant degradation on two others, leading to the lowest overall average score. In contrast, the foundation models demonstrated more consistent cross-dataset performance. Therefore, this result highlighting a critical issue: SFT enhances in-distribution performance but weakens generalization to unseen scenarios. (e.g., novel tools, invocation formats).
@@ 4.1 Training Data Preparation
The following data are utilized in RL training. (1) ToolACE (liu2024toolace): It is a general-purpose tool-use dataset, where the model learns when to invoke tools and when to respond directly, thereby enhancing decision-making in multi-step interactions. (2) xLAM (liu2024apigen): This is a compositional dataset that requires one or multiple tool calls per turn. It encourages the model to reason about tool dependencies and actively plan diverse tool-calling actions. We also include an irrelevance-augmented subset [^1] https://huggingface.co/datasets/MadeAgents/xlam-irrelevance-7.5k originating from xLAM.
@@
Data Mask. Recent studies (lin2024hammer; chen2025acebench) have shown that naming preferences in tool descriptionss can significantly degrade model robustness when testing environments diverge from training conventions. To mitigate this issue, we adopt a name-masking strategy aligned with lin2024hammer, which masks function names (e.g., calculate_sumfunc_1) and parameter names (e.g., input_listparam_1). It redirects the model’s attention to tool descriptions and argument semantics, reduces overfitting to superficial naming patterns, thereby improving tool-agnostic reasoning in open-domain settings.
@@ 4.2 Generalization Guided Reward Design
To enhance tool generalization within the GRPO framework, we introduce a dynamic generalization-guided reward design that combines flexible exploration with structured convergence. Building on prior rule-based reward mechanisms (qian2025toolrl; li2025torl; jin2025search), our formulation decomposes the total reward $\mathcal{R}_{\text{final}}$ into two components:
@@
To further enhance the model’s tool-integrated reasoning in context, we incorporate two tool-specific feedback signals into the strict reward $\mathcal{R}_{\text{strict}}$: (1) multi-tool collaboration: +0.3+0.3 reward for reinforcing correct collaborative tool usage patterns in multi-step tool chains or dialogs. (2) parameter value error: +0.3+0.3 penalty per invalid parameter value to enforce exploring precise context grounding because it demands high-order reasoning from context (zeng2025boosting; lin2024hammer).
@@ 4.3 RL Training with Generalization-guided Reward
To train LLMs from Zero model through scaling reinforcement learning, we employ Group Relative Policy Optimization (GRPO) (deepseekai2025r1; shao2024deepseekmath), that unlocks the model’s reasoning capabilities at test time.
@@
here, $\epsilon$ is a hyperparameter. Following yu2025dapo, we remove the KL divergence regularization from the GRPO objective. And $A_{i}$ is the advantage computed using a group of rewards $\{r_{1},r_{2},\dots,r_{G}\}$ corresponding to the completions within each group:
@@ 5.1 Experimental Setup
Evaluation Dataset. The BFCL evaluates the LLMs ability to invoke functions in the real-world by actually triggering the API call and comparing responses, provides a comprehensive dataset comprising 4k+ instances (updating), consisting of Non-live, Live (with user-contributed complex tools avoiding contamination), Multi-turn subset. Other benchmarks, namely API-Bank (li2023api), Nexus Raven (srinivasan2023nexusraven), Tool-Alpaca (tang2023toolalpaca), and Seal-Tools (wu2024seal), are elaborated in Appendix B.
Baselines (1) Vanilla Model: the original model without additional training (e.g., Llama3.1-series, Qwen2.5-series). (2) SFT Models: instruct models fine-tuned on supervised data, to assess whether GRPO training outperforms standard SFT, including ToolACE-8B (trained in ToolACE), xLAM-series (trained in xLAM)(zhang2024xlam), and Hammer-series (trained on xLAM with function mask to enhance generalization) (lin2024hammer). (3) API-based closed-source models (e.g., GPT-series, Gemini-series). (4) R1-like Model: models trained using GRPO with SFT as the RL paradigm, such as QwQ-32B (qwq32b), Tool-N1 series (single turn tool-use models trained in mixed ToolACE and xLAM data) (zhang2025nemotron), and ToolRL(trained in subset of mixed ToolACE and xLAM data) (qian2025toolrl). Instead, our Tool-Zero series trained with pure RL without SFT (i.e., R1-Zero).
@@ 5.3.1 Ablation Study
Additionally, we conduct a hyperparameter ablation study on the progressive reward strategy by varying two key parameters: transition midpoint $t_{m}\in\{0,25,50,100\}$ and steepness factor $\kappa\in\{0.2,1.0\}$ (controlling reward transition slope). Results (Figure 5) show that a smaller transition midpoint ($t_{m}=25$) yields the best performance, while larger values ($t_{m}\geq 50$) lead to degradation. This aligns with prior observations (pan2024feedback; zhang2025nemotron) that excessive exploration in fine-grained schemes may induce reward hacking and overfitting to superficial cues. Also, a lower steepness factor consistently outperforms, indicating that gradual reward shaping stabilizes training. These findings validate the design choices in GG-GRPO’s progressive reward mechanism.
@@ A.1 The Implementation Settings
The BFCL is an evolving benchmark. For our study, we utilized the version checked out on February 26, 2024. Other benchmarks include: (1) API-Bank with 314 tool-use dialogues and 753 API calls, evaluating known API invocation (L-1) and candidate list retrieval/calling (L-2), we report their average result in evaluation; (2) Nexus Raven API Evaluation offering 318 test examples across 65 APIs for function-calling assessment; (3) Tool-Alpaca’s 271 synthetic tool-use instances in 50 categories (100 simulated tests used); (4) Seal-Tools, a recent benchmark with 4,076 auto-generated APIs across life domains. The BFCL assesses models using Abstract Syntax Tree Evaluation and Executable Function Evaluation Accuracy, and the other benchmarks assesses models using Function and Parameter matching F1 score (lin2024hammer). For the evaluation, each model was administered the same system prompt as its training counterpart.
@@ Appendix C System Thinking Template
We adopt a lightweight prompting schema to elicit tool-use capabilities from the LLM, drawing inspiration from prior work deepseekai2025r1; openr1. As illustrated in Figure 6, the template explicitly instructs the model to encapsulate intermediate reasoning within <think>...</think> tags, followed by the final answer enclosed in <answer>...</answer> tags. By allowing the model greater freedom in articulating its reasoning process, we aim to enhance generalization across diverse tool integration scenarios. Additionally, this design facilitates seamless adaptation to complex tool-augmented reasoning tasks.
@@ Appendix D Model Scale Analysis
Additionally, for the experiment on algorithmic effectiveness, the objective was to evaluate the reward function on native PPO (song2025r1; zeng2025simplerl) (i.e., GG-PPO) and GRPO algorithms (i.e., GG-GRPO), as shown in Table 9. The results revealed that our reward module with GRPO outperformed PPO, demonstrating robust and consistent gains. This highlights GRPO’s superior adaptability to the reward framework for tool learning, which can be attributed to the reward function being designed for rule-based rewards in algorithms like GRPO (unlike PPO, which is natively for model-based rewards) and recent studies showing GRPO’s advantage in achieving "aha moments" compared to PPO-family algorithms (qian2025toolrl; deepseekai2025r1; muennighoff2025s1).